OpenStudy (anonymous):

A bacteria culture initially contains 3000 bacteria and doubles every half hour. Find the size of the baterial population after 100 minutes. ? Find the size of the baterial population after 4 hours?

OpenStudy (anonymous):

A)100/30=3 1/3 * 3000 B) 8x3000

OpenStudy (anonymous):

did that help???

OpenStudy (anonymous):

yes it did thank you :)!!!:D

OpenStudy (anonymous):

NP

OpenStudy (l):

B(t) = B(0) * 2^t B(0) = 3000 After 100 mins: B = B(0) * 2^(100/30) ~ 30238,1052 -> B = 30238 After 4 hours: B = B(0) * 2 ^(4/0.5) = 3000 * 2^8 = 3000 * 64 = 192000 -> B = 192000 *